OpenStudy (anonymous):

An 8-pound bag of tomatoes costs $5.60. At the same rate, what would be the cost of 11 pounds of tomatoes

OpenStudy (anonymous):

$7.70

OpenStudy (hoblos):

5.60/8 = x/11 x= 11*5.6/8=$7.70